what is the slope of the line perpendicular to
y=2/5x+7

Answer by fcabanski(1391)

y=mx+b is the equation of a line where m = slope. The slope of the given line is 2/5.


The slope of a line perpendicular to line a is the negative reciprocal of the slope of line a. So the slope of a line perpendicular to this line, with slope 2/5, is the negative reciprocal of 2/5 = -5/2.


Contact me for private tutoring via online voice.